# Minimising surface-code failures using a color-code decoder

### v2

@@ 5 Surface code performance under depolarising noise
In this section, we study the surface code under depolarising Pauli noise on data qubits and perfect measurements. We compare the performance of the surface-code decoder we have introduced to the conventional matching decoder. Under the mapping, these correspond to the unified decoder and the restricted decoder, respectively.
@@ 5.2 Path-counting of failure mechanisms
We now evaluate the number of least-weight errors that lead the surface code to failure using the restricted-lattice decoder. This decoder cannot distinguish between Pauli-Z or Pauli-Y errors. As such we count the number of weight $d/2$ configurations of errors along a major row of the color-code lattice, where we say a major row is a row of $d$ red plaquettes. In contrast, a minor row of the lattice is a row of $d-1$ red color-code plaquettes. Since the lattice is square, it supports $d$ major rows and $d-1$ minor rows.
@@ 5.4 Thresholds
where $x$ is the re-scaled error rate $x=(p-p_{\textrm{th}})^{1/\nu}$ [28]. We find a fitted threshold of $p_{\textrm{th}}\leavevmode\nobreak\ \sim 15.4(7)\%$ for the restricted decoder under depolarising errors using the data shown in Fig. 9, where we obtain a fitted critical exponent $\nu\approx 0.611$, and Taylor fit constants $A=0.471$, $B=0.540$ and $C=0.153$. For the unified decoder, we find that the threshold estimate varies slightly with the chosen values of $w_{A}$ and $w_{B}$. We considered multiple weights $w_{B}$ with fixed $w_{A}=1$ for edges of the matching graph in Fig. 4(c), and find a threshold of $p_{\textrm{th}}\leavevmode\nobreak\ \sim 15.2(1)\%$ using $w_{A}=1$ for $w_{B}=0.5$, as shown in Fig. 9 where $\nu=0.664$, $A=0.606$, $B=0.452$ and $C=0.127$. We provide the data for different weight configurations in Appendix A, where the threshold is still comparable to that of the standard surface code, with a minor discrepancy.
@@ 6.2 Low error-rate performance
where $n$ is the number of physical qubits in the color code, $|E|$ the weight of an error operator $E$. Once again, we adopt the sequence $p_{j+1}=p_{j}2^{\pm 1/\sqrt{w_{j}}}$ to split ratios [63] to minimise statistical error, where $w_{j}=\text{max}(d/2,p_{j}n)$ for a code distance $d$. We numerically interpolate between initial estimates obtained using Monte Carlo sampling at physical error rate $p=5\%$, and the analytical path counting results in the low error rate regime shown in Sec. 6.1. The results are shown in Fig. 10 where a MWPM decoder was executed on the color code restricted and unified lattices respectively, under bit-flip noise and perfect measurements. For the distances considered, the analytic estimates in Eqs. 20 and 22 accurately predict the dominant behaviour of the error chains in the low error rate regime, as indicated by the solid line fits. We note that in order to correct for all expected minimum-weight configurations for bit-flip errors on the color code there is a constraint on the edge weights, namely that $w_{A}>w_{B}$, which does not exist is the case of depolarising errors on the even-distance surface code.
@@ 6.3 Thresholds
The restricted decoder recovers the threshold $p_{\textrm{th}}\leavevmode\nobreak\ \sim 10.2\%$ under independent and identically distributed bit-flip noise, see Fig. 11, where we obtain a fitted critical exponent $\nu=0.628$, and Taylor fit constants $A=-0.0419$, $B=1.136$ and $C=0.165$ for Eq. 18. This is consistent with that of the surface code on a square lattice [28, 45]. In the case of the unified decoder, we considered multiple weights $w_{A}$ with $w_{B}=1$ for edges of the matching graph in Fig. 4(c), and find a peak in threshold at $p_{\textrm{th}}\sim 10.1(3)\%$ around $w_{A}=2.1$ (Fig. 11). Surprisingly, the unified decoder demonstrates a marginally lower threshold than the restricted decoder.
@@ 7 Discussion
It will be interesting to determine how the performance of our decoder compares to other practical decoders that are designed to account for depolarising noise, both in terms of error-correction performance and runtime. For instance, both iterative [45, 56] and belief-matching decoders [57] have been proposed to augment matching decoders with the ability to take advantage of correlations between the syndromes measured by the star and plaquette operators. To complete such a comparison, it will first be necessary to determine the performance of other decoding algorithms in the limit of low error rate. Of course, the matching decoder we have presented here is compatible with other decoding techniques such as belief-matching. It may be interesting to design versatile decoders for other codes by adapting these techniques. One could also readily combine our decoder with fast decoders such as union-find [14, 72, 73] or weighted union-find [74] to improve its runtime. Of course, we can also generalise our versatile matching decoder to deal with non-trivial syndrome correlations that occur in fault-tolerant error correction [28, 50, 17], and may be adapated for use in other variations of the surface code with boundaries such as Floquet codes [75, 54]. One might also explore generalising our decoding method to surface codes with alternative topologies, or perhaps using other color code symmetries. Both of these directions are discussed in Appendix F of Ref. [48].
Broadly speaking, we have shown that we can obtain a matching decoder for the surface code capable of correcting depolarising noise by mapping its syndrome onto an enlarged state space, namely, the syndrome of the color code. It will be interesting to explore this method further to determine the extent to which it generalises. For instance, it is compelling to investigate whether we obtain near optimal performance using our mapped decoder for different lattice geometries [66]. Indeed, in preliminary work we have already identified subtle differences in the decoding problem if we consider odd distance surface codes. We will report on these differences in a updated version of this present manuscript. More generally, it would be exciting to consider other decoding problems on extended Hilbert spaces. This may lead to practical, high performance decoders for more exotic codes. We leave this exploration to future work.
@@
Appendix A Thresholds for the even-distance surface code using varied weights

### v3

@@ 5 Surface code performance under depolarising noise
In this section, we study the surface code under depolarising Pauli noise on data qubits and perfect measurements. We compare the performance of the surface code decoder we have introduced to the conventional matching decoder. Under the mapping, these correspond to the unified decoder and the restricted decoder, respectively.
@@ 5.2 Path-counting of failure mechanisms
We now evaluate the number of least-weight errors that lead the surface code to failure using the restricted-lattice decoder. This decoder cannot distinguish between Pauli-Z or Pauli-Y errors. As such we count the number of weight $d/2$ configurations of errors along a major row of the color code lattice, where we say a major row is a row of $d$ red plaquettes. In contrast, a minor row of the lattice is a row of $d-1$ red color code plaquettes. Since the lattice is square, it supports $d$ major rows and $d-1$ minor rows.
@@ 5.4 Thresholds
where $x$ is the re-scaled error rate $x=(p-p_{\textrm{th}})^{1/\nu}$ [28]. We find a fitted threshold of $p_{\textrm{th}}\leavevmode\nobreak\ \sim 15.4(7)\%$ for the restricted decoder under depolarising errors using the data shown in Fig. 9, where we obtain a fitted critical exponent $\nu\approx 0.611$, and Taylor fit constants $A=0.471$, $B=0.540$ and $C=0.153$. For the unified decoder, we find that the threshold estimate varies slightly with the chosen values of $w_{A}$ and $w_{B}$. We considered multiple weights $w_{B}$ with fixed $w_{A}=1$ for edges of the matching graph in Fig. 4(c), and find a threshold of $p_{\textrm{th}}\leavevmode\nobreak\ \sim 15.2(1)\%$ using $w_{A}=1$ for $w_{B}=0.5$, as shown in Fig. 9 where $\nu=0.664$, $A=0.606$, $B=0.452$ and $C=0.127$. We provide the data for different weight configurations in Appendix B, where the threshold is still comparable to that of the standard surface code, with a minor discrepancy.
@@ 6.2 Low error-rate performance
where $n$ is the number of physical qubits in the color code, $|E|$ the weight of an error operator $E$. Once again, we adopt the sequence $p_{j+1}=p_{j}2^{\pm 1/\sqrt{w_{j}}}$ to split ratios [63] to minimise statistical error, where $w_{j}=\text{max}(d/2,p_{j}n)$ for a code distance $d$. We numerically interpolate between initial estimates obtained using Monte Carlo sampling at physical error rate $p=5\%$, and the analytical path counting results in the low error rate regime shown in Sec. 6.1. The results are shown in Fig. 10 where a MWPM decoder was executed on the color code restricted and unified lattices respectively, under bit-flip noise and perfect measurements. For the distances considered, the analytic estimates in Eqs. 20 and 22 accurately predict the dominant behaviour of the error chains in the low error rate regime, as indicated by the solid line fits. We note that in order to correct for all expected minimum-weight configurations for bit-flip errors on the color code there is a constraint on the edge weights, namely that $w_{A}>w_{B}$, which does not exist in the case of depolarising errors on the even-distance surface code.
@@ 6.3 Thresholds
The restricted decoder recovers the threshold $p_{\textrm{th}}\leavevmode\nobreak\ \sim 10.2\%$ under independent and identically distributed bit-flip noise, see Fig. 11, where we obtain a fitted critical exponent $\nu=0.628$, and Taylor fit constants $A=-0.0419$, $B=1.136$ and $C=0.165$ for Eq. 18. This is consistent with that of the surface code on a square lattice [28, 45]. In the case of the unified decoder, we considered multiple weights $w_{A}$ with $w_{B}=1$ for edges of the matching graph in Fig. 4(c), and find a peak in threshold at $p_{\textrm{th}}\sim 10.1(3)\%$ around $w_{A}=2.1$ (Fig. 11). Surprisingly, the unified decoder demonstrates a marginally lower threshold than the restricted decoder. This pattern is consistent with the threshold comparison applied to the surface code in Sec. 5.4. We hypothesize that in the near-threshold regime, the presence of additional defects on the unified lattice contributes to an increased number of incorrect minimum-weight paths that the decoder may choose, leading to this reduction.
@@ 7 Discussion
It will be interesting to determine how the performance of our decoder compares to other practical decoders that are designed to account for depolarising noise, both in terms of error-correction performance and runtime. For instance, both iterative [45, 56] and belief-matching decoders [57] have been proposed to augment matching decoders with the ability to take advantage of correlations between the syndromes measured by the star and plaquette operators. To complete such a comparison, it will first be necessary to determine the performance of other decoding algorithms in the limit of low error rate. Of course, the matching decoder we have presented here is compatible with other decoding techniques such as belief-matching and two-stage decoding [65, 74], at the cost of added computational complexity, and it may be interesting to design versatile decoders for other codes by adapting these techniques. However, it is important to note that our decoder only requires a single invocation of an MWPM decoder to address Pauli correlations. As a result, it benefits from the expected linear time complexity of MWPM [31], with only a constant factor increase of 4 in the size of the matching graph. One could also readily combine our decoder with fast decoders such as union-find [14, 75, 76] or weighted union-find [77] to improve its runtime. We note that we can also generalise our versatile matching decoder to deal with non-trivial syndrome correlations that occur in fault-tolerant error correction [28, 50, 17], and may be adapted for use in other variations of the surface code with boundaries such as Floquet codes [78, 54]. One might also explore generalising our decoding method to surface codes with alternative topologies, or perhaps using other color code symmetries. Both of these directions are discussed in Appendix F of Ref. [48].
Broadly speaking, we have shown that we can obtain a matching decoder for the surface code capable of correcting depolarising noise by mapping its syndrome onto an enlarged state space, namely, the syndrome of the color code. It will be interesting to explore this method further to determine the extent to which it generalises. For instance, it is compelling to investigate whether we obtain near-optimal performance using our mapped decoder for different lattice geometries [66]. Another important case to consider is the circuit noise model, where decoding is conducted on a $(2+1)$-dimensional syndrome history. Trivially, we can adapt the matching graph obtained from the unified lattice for this setting by extending it along the time direction in the standard way [7, 41, 55]. Nevertheless, it will be very interesting if the novel decoding methods we have proposed can be generalised to offer an advantage in correcting the types of correlated errors the circuit noise model can introduce. Perhaps, for instance, the edge weights on the unified lattice can be chosen appropriately to account for correlated errors such as hook errors. Perhaps there is even some other alternative lattice that can be found to offer an advantage in dealing with this more complicated noise model. More generally, it is exciting to consider other decoding problems on extended Hilbert spaces. This may lead to practical, high-performance decoders for more sophisticated codes, as well as more exotic correlated noise models. We leave this exploration to future work.
@@
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request. The source code and data used in the figures in this manuscript can be accessed at https://github.com/abenhemou/unimatch.
Appendix A Pseudocode for unified decoding on the surface code
Here, we outline the Unified decoder presented in the main text. The procedure requires a set of syndrome measurements of the color code stabilizers $\sigma_{\textrm{c.c}}$, a subroutine which restricts $\sigma_{\textrm{c.c}}$ to syndromes on restricted lattices, and a subroutine which builds the unified lattice as prescribed in Fig. 4. A minimum-weight perfect matching decoder mwpmdecoder is then used to decode the syndrome and provide a correction $\bar{C}$. The correction operator returns the code to the codespace, and its success is given by the condition that the decoding algorithm has determined that the parity of the edges in the matching solution crossing a chosen logical representative $\bar{Z}$ is equal to the parity of Pauli errors which occurred on qubits $q\in\bar{Z}$, as mentioned in Sec. 3.2.
1
2Require:
Restricted lattices $\mathcal{R}_{\mathbf{u}}=\{\{S_{f}\}_{\mathbf{col}(f)\neq\mathbf{u}}\}$ on a manifold $G_{\mathcal{R}_{\mathbf{u}}}$
Unified lattice constructor $\mathcal{U}(.)$
A MWPM decoder mwpmdecoder
3Input: A color code syndrome $\sigma_{\textrm{c.c.}}$ on the color code manifold $G_{\textrm{c.c}}$
4Output: A correction operator $\bar{C}$
/* Map syndromes to restricted lattices */
5
6foreach $\textbf{{u}}\in\{\textbf{{r}},\textbf{{g}},\textbf{{b}}\}$ do
7
8      foreach $\sigma_{i}\in\sigma_{\textrm{c.c.}}$ do
9             if $color(\sigma_{i})=\textbf{{u}}$ then
10
11                  $\mathcal{R}_{\mathbf{u}}(\sigma_{i})\text{ on }G_{\mathcal{R}_{\mathbf{u}}}\leftarrow\sigma_{i}\text{ on }G_{\textrm{c.c.}}$
12             end if
13
14       end foreach
15
16 end foreach
17
/* Construct unified lattice, and map syndrome on the unified lattice */
18
19$(G_{\mathcal{U}},\sigma_{i,\mathcal{U}})\leftarrow\mathcal{U}\left(G_{\mathcal{R}_{\mathbf{r}}},G_{\mathcal{R}_{\mathbf{g}}},G_{\mathcal{R}_{\mathbf{b}}},\mathcal{R}_{\mathbf{r}}(\sigma_{i}),\mathcal{R}_{\mathbf{g}}(\sigma_{i}),\mathcal{R}_{\mathbf{b}}(\sigma_{i})\right)$
/* Decode */
20 $\bar{C}=\texttt{mwpmdecoder}(G_{\mathcal{U}},\sigma_{i,\mathcal{U}})$
/* Apply Pauli-$Z$ corrections to qubits along the shortest paths determined by the decoder */
21
22Apply $\prod_{q\in\bar{C}}Z_{q}$
Algorithm 1 Unified Decoder
Appendix B Thresholds for the even-distance surface code using varied weights